OpenStudy (anonymous):

Last week, Eva spent 8 and one-eighth hours on homework. This week, she spent eleven and seven-eighths hours. How many more hours did she spend this week? 2 3/4 2 1/4 3 3/4 3 1/4

OpenStudy (texaschic101):

11 7/8 - 8 1/8 Lets set aside the whole numbers for now. 7/8 - 1/8 = 6/8 reduces to 3/4 Now bring back the whole numbers... 11 - 8 = 3 combine and you get 3 3/4

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

nope thanks

OpenStudy (texaschic101):

sure thing :)